40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Aug 24 00:39:47 2017, gautam, Update, LSC, DRMI locking attempt 
    Reply  Fri Aug 25 01:20:52 2017, gautam, Update, LSC, DRMI locking attempt 
       Reply  Wed Aug 30 19:49:33 2017, gautam, Update, LSC, REFL55 demod board debugging REFL55_demod_check.pdf
          Reply  Thu Aug 31 00:52:52 2017, gautam, Update, LSC, REFL55 whitening board debugging REFL55_whtCheck.pdfREFL55_whtChk.tar.gz
             Reply  Thu Aug 31 03:31:15 2017, gautam, Update, LSC, DRMI re-locked! DRMI_1f.pngDRMI_relocked.pdf
                Reply  Mon Sep 4 16:30:06 2017, gautam, Update, LSC, Oplev loop tweaking 
                   Reply  Mon Sep 4 18:18:29 2017, rana, Update, LSC, dewhite switching: FOTON settings 
                   Reply  Tue Sep 5 02:07:49 2017, gautam, Update, LSC, Low Noise DRMI attempt MICH_err_comp.pdfdeWhitenedCoil.pdf
                      Reply  Wed Sep 6 23:06:30 2017, gautam, Update, LSC, Coil de-whitening switching investigation BScoils.pdfITMcoils.pdf
                         Reply  Fri Sep 15 16:00:33 2017, gautam, Update, LSC, Sensing measurement DRMI1f_Sep5.pdf
                         Reply  Fri Sep 15 17:08:58 2017, gautam, Update, LSC, Coil de-whitening switching investigation DRMI_IntensityNoise.pdf
                            Reply  Sat Sep 16 10:56:19 2017, rana, Update, LSC, Coil de-whitening switching investigation 
                               Reply  Fri Sep 22 18:12:27 2017, gautam, Update, LSC, DAC noise measurement (again) 
                                  Reply  Wed Sep 27 22:25:21 2017, gautam, Update, LSC, DAC noise measurement (again) coilNoises.pdf
                                     Reply  Thu Sep 28 06:35:44 2017, Chris, Update, LSC, DAC noise measurement (again) 
                                        Reply  Thu Sep 28 23:50:04 2017, gautam, Update, LSC, DAC noise measurement (again) 
                                           Reply  Fri Sep 29 18:36:25 2017, gautam, Update, LSC, DAC noise measurement (again) 
                               Reply  Tue Oct 3 01:32:39 2017, gautam, Update, LSC, Laser intensity noise coupling to MICH (simulated) MICH_intensityNoiseCoupling.pdfMICH_intensityNoiseASD.pdf
Message ID: 13281     Entry time: Thu Aug 31 03:31:15 2017     In reply to: 13280     Reply to this: 13289
Author: gautam 
Type: Update 
Category: LSC 
Subject: DRMI re-locked! 

After our Demod/Whitening electronics investigations suggested nothing obviously wrong, I decided to give DRMI locking another go tonight.

Surprisingly, there was no evidence of REFL55 behaving weirdly tonight, and I was able to easily lock the DRMI on 1f error signals using the recipe I've been using in the last few months.

Not sure what to make of all this frown.

I got in a ~15 minute lock, but I wasn't prepared to do any sort of characterization/ sensing / attempt to turn on coil-dewhitening, and I'm too tired to try again tonight. I was however able to whiten the error signals, as I have been able to do in the past. There is a ~45Hz bump in MICH that I haven't seen in the past.

I'll try and do some characterization tomorrow eve, but it's encouraging to at least get back to the pre-FB-failure state of locking.

Attachment 1: DRMI_1f.png  52 kB  Uploaded Thu Aug 31 04:32:00 2017  | Hide | Hide all
DRMI_1f.png
Attachment 2: DRMI_relocked.pdf  52 kB  Uploaded Thu Aug 31 04:33:10 2017  | Hide | Hide all
DRMI_relocked.pdf
ELOG V3.1.3-